1.   A tank initially contains 50 gal of fresh water. At t=0, brine solution containing 2 lb of salt per gallon is poured into the tank at the rate of 12 gal/min, while the well stirred mixture leaves the tank at the rate of 8 gal/min. what is the amount of salt at the end of 5 minutes? How long will it take to obtain an amount of 50 lb? ans. 88.98 lb; 2.47 min
  2. Given an RC series circuit that has an emf source of 50 volts, a resistance of 20 kilo ohms, a capacitance of 6 microfarad and the initial charge of the capacitor of the capacitor is 1 micro coulomb. What is the charge in the capacitor at the end of 0.01 second? What is the current in the circuit at the end of 0.05 seconds? Ans. 24.91µC; 1.64mA
1
Expert's answer
2021-11-17T15:24:16-0500

1. Let "s(t) =" amount, in lb of salt at time "t." Then we have

"\dfrac{ds}{dt}="(rate of salt into tank) − (rate of salt out of tank)


"\dfrac{ds}{dt}=2\cdot12-\dfrac{8s}{50+(12-8)t}"

So we get the differential equation


"\dfrac{ds}{dt}=24-\dfrac{8s}{50+4t}"

"\dfrac{ds}{dt}+\dfrac{4s}{25+2t}=24"

Integrating factor


"\mu(t)=e^{\int 4dt/(25+2t)}=(25+2t)^{2}"

"\dfrac{d}{dt}((25+2t)^{2}s)=24(25+2t)^{2}"

"\int d((25+2t)^{2}s)=24\int(25+2t)^{2}dt"

"(25+2t)^{2}s=24(\dfrac{1}{6})(25+2t)^{3}+c_1"

"s(t)=4(25+2t)+c_1(25+2t)^{-2}"

"s(0)=100+c_1(25)^{-2}=0"

"c_1=-62500"

"s(t)=4(25+2t)-62500(25+2t)^{-2}"

"s(5)=4(25+2(5))-62500(25+2(5))^{-2}"

"s(5)=88.98\ lb"


"s(t)=4(25+2t)-62500(25+2t)^{-2}=50"

Let "y(t)=s(t)-50"


"y(t)=4(25+2t)-62500(25+2t)^{-2}-50"

Solve graphically the equation "y=0"


"t=2.47\ min"


2.


"Rq'+\dfrac{1}{C}q=V"

"2\cdot10^{4}q'+\dfrac{1}{6\cdot10^{-6}}q=50"

"q'+\dfrac{25}{3}q=0.0025"

"q'=-\dfrac{25}{3}(q-0.0006)"

"\dfrac{dq}{q-0.0003}=-\dfrac{25}{3}dt"

Integrate


"q=0.0003+c_1e^{-25t\/3}"

Given "q(0)=10^{-6}"


"0.000001=0.0003+c_1"

"c_1=-0.000299"

"q(t)=0.0003-0.000299e^{-25t\/3}"

"q(0.01)=0.0003-0.000299e^{-25(0.01)\/3}"

"=24.91\cdot10^{-6} (C)"

"24.91\ \mu C"


"i(t)=q'(t)=-0.000299(-\dfrac{25}{3})e^{-25t/3}"

"=\dfrac{0.007475}{3}e^{-25t/3}"

"i(0.05)=\dfrac{0.007475}{3}e^{-25(0.05)/3}=0.0016426(A)"

"=1.64(\ mA)"

"1.64\ mA"